You are traveling 6 mi. east and 8 mi. north to Franklin Lane. The speed limit is 40 mph. How many minutes will it take to get there in a 14 mi stretch ?
Answer by rwm(914)

40x=14
x=14/40
x=3.5/10=.35 hr
x=.35*60=21 minutes